Evaluate the integral of e^(5x)dx


1
Expert's answer
2021-02-24T12:47:26-0500

I = "\int" e5x "dx"


We know that,

"\int" e(ax) "dx" = "\dfrac{e^{ax}}{\dfrac{d}{dx}(ax)}" + C = "\dfrac{e^{ax}}{a}" + C

On comparing this formula with our question, we find that

a = 5


So, I = "\int" e5x "dx" = "\dfrac{e^{5x}}{5}" + C